In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all. How many votes did each candidate receive??
Answer by ptaylor(2198)

Lex x=number of votes for the loser
Then x+220=number of votes for the winner
loser votes + winner votes = 810, so:
x+x+220=810 subtract 220 from both sides
x+x+220-220=810-220 collect like terms
2x=590 divide both sides by 2
x=295-------------------------number of votes for the loser
x+220=295+220=515------------------------number of votes for the winner
CK
295+515=810
810=810
Hope this helps----ptaylor